find the solution and ususal about the solution uxx=1/100ut , u(0,t)=Ae-dt u(x,0)=0 and ux approaches to 0 as x approahes to zero



1
Expert's answer
2021-01-27T17:14:58-0500

"\displaystyle
u(x, t) = X(x)T(t)\\

u_{xx}(x, t) = X''(x)T(t)\\

u_{t} = X(x)T'(t)\\

X''(x)T(t) = \frac{1}{100}X(x)T'(t)\\

\frac{100X''(x)}{X(x)} = \frac{T'(t)}{T(t)} = K\\

\int \frac{T'(t)}{T(t)}\,\mathrm{d}t = \int K \,\mathrm{d}t\\

\ln\left(T(t)\right) = Kt + C\\

T(t) = e^{Kt + C} = Ae^{Kt}\\

\frac{100X''(x)}{X(x)} = K\\

\frac{X''(x)}{X(x)} = \frac{K}{100}\\

X''(x) - \frac{KX(x)}{100} = 0\\

\textsf{If}\,\, K\,\, \textsf{is chosen to be negative}\\

\textsf{say}\,\, K = -\lambda^2,\,\,\textsf{then}\,\,\\

X'' = -\frac{\lambda^2X(x)}{100}\\

\therefore X = B\cos\left(\frac{x}{10}\right) + C\sin\left(\frac{x}{10}\right)\\


\begin{aligned}
u(x, t) &= Ae^{-\lambda^2t}\left(B\cos\left(\frac{x}{10}\right) + C\sin\left(\frac{x}{10}\right)\right)\\
&= e^{-\lambda^2t}\left(D\cos\left(\frac{x}{10}\right) + E\sin\left(\frac{x}{10}\right)\right)
\end{aligned} \\

u(0, t) = De^{-\lambda^2t}= Ae^{-dt} \\

\therefore A = D,\,\, \textsf{and}\,\, d = \lambda^2\\

u(x, 0) = D\cos\left(\frac{x}{10}\right) + E\sin\left(\frac{x}{10}\right) = 0 \\


\lim_{x \to 0} u(x, 0) = \lim_{x \to 0} \left(D\cos\left(\frac{x}{10}\right) + E\sin\left(\frac{x}{10}\right)\right) = 0 \\

\implies D = 0.\\

\begin{aligned}
\therefore u(x, t) &= Ee^{-dt}\sin\left(\frac{x}{10}\right)
\end{aligned}"